OpenStudy (anonymous):

Which answer shows this equation written in standard form? 9 – 4(x – y) = –6x + 6

OpenStudy (anonymous):

A. 2x + 4y = 3 B. –10x + 4y = –3 C. 2x – 4y = 3 D. 2x + 4y = –3

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

youre welcome(:

OpenStudy (anonymous):

thanks for the medal!!

OpenStudy (anonymous):

no prob :)